Find the electric force between two protons separated by a distance of 1 fermi (1 fermi = 10-15 m). The protons in a nucleus remain at a separation of this order.

Grade:10

3 Answers

Jitender Pal
9 years ago
Sol. Charge on each proton = a= 1.6 10^-19 coulomb Distance between charges = 10 * 10^-15 metre = r Force = kq^2/r^2 = 9 * 10^9 * 1.6 * 1.6 * 10^-38/10^-38 = 9 * 2.56 * 10 = 230.4 Netwon
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Apoorva Arora IIT Roorkee
9 years ago
force is given by
F=\frac{kq_{1}q_{2}}{r^{2}}
putting all the values
F= 230.4 newtons
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty